# Carlo Orientale Caputo SISSA - International School for Advanced Studies, 34136 Trieste, Italy Elias Seiffert

### v3

@@ 1.1 The relevance
We describe representations $p(\mathbf{s})$ in terms of the variable $E_{\mathbf{s}}=-\log_{2}p(\mathbf{s})$, which is the minimal number of bits needed to represent state $\mathbf{s}$. The average coding cost $H[\mathbf{s}]=\mathbb{E}\left[E_{\mathbf{s}}\right]$ is the usual Shannon entropy and counts the number of bits needed to describe one point of the dataset. Following Ref. [36], we shall call $H[\mathbf{s}]$ the resolution.
@@
where $p(E)=W(E)2^{-E}$ is the probability that a state $\mathbf{s}$ randomly drawn from $p(\mathbf{s})$ has $E_{\mathbf{s}}=E$, and $W(E)$ is the number of states $\mathbf{s}$ with $E_{\mathbf{s}}=E$. The principle of maximal relevance [36] postulates that maximally informative representations should achieve a maximal value of $H[E]$, which correspond to a uniform distribution of coding costs ($p(E)={\rm const}$ or $W(E)=W_{0}2^{E}$). Representations where coding costs are distributed uniformly should be promoted for the reason that, in an optimal representation, the number $W(E)$ of states $\mathbf{s}$ that require $E$ bits to be represented should match as closely as possible the number ($2^{E}$) of codewords that can be described by $E$ bits. Representation of maximal relevance with a given resolution also have an exponential degeneracy of states $W(E)=W_{0}e^{\nu E}$, with $\nu$ that depends on $H[\mathbf{s}]$. Note that states $\mathbf{s}$ and $\mathbf{s}^{\prime}$ with very different coding costs $E_{\mathbf{s}}$ and $E_{\mathbf{s}^{\prime}}$ can be distinguished by their statistics, because they would naturally belong to different typical sets [^7] By the law of large numbers, typical samples of weakly interacting variables all have approximately the same coding cost, a fact knowns as the asymptotic equipartition property [40]. Following Ref. [41], we take the view that a trained DBN distinguishes the points in a dataset in different typical sets.. Representations that maximise the relevance harvest this benefit in discrimination ability that is accorded by statistics alone.
1.2 The Hierarchical Feature Model
The HFM encodes the principle of maximal relevance. It describes the distribution of a string $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ of binary variables that we take as indicators of whether each of $n$ features is present ($s_{i}=1$) or not ($s_{i}=0$). Features are organised in a hierarchical scale of detail and we require that the occurrence of a feature $s_{k}=1$ at level $k$ does not provide any information on whether lower order features are present or not. This means that, conditional on $s_{k}=1$, all lower order features are as random as possible, i.e. $H[\mathbf{s}_{1:k-1}|s_{k}=1]=k-1$ in bits. This requirement implies that the Hamiltonian $E_{\mathbf{s}}$ should be a function of $m_{\mathbf{s}}=\max\{k:~{}s_{k}=1\}$, with $m_{\mathbf{s}}=0$ if $\mathbf{s}=\mathbf{0}$ is the featureless state [^8] This is because $p(\mathbf{s}|m_{\mathbf{s}}=k)=2^{-k+1}$ so that $p(\mathbf{s})=p(\mathbf{s}|m_{\mathbf{s}})p(m_{\mathbf{s}})=p(m_{\mathbf{s}})/2^{m_{\mathbf{s}}-1}$. ($s_{i}=0~{}\forall i$). Since there are $2^{k-1}$ states with $m_{\mathbf{s}}=k$, the principle of maximal relevance (i.e. the requirement that $W(E)=W_{0}e^{\nu E}$) excludes all functional forms between $E_{\mathbf{s}}$ and $m_{\mathbf{s}}$ that are not linear. This leads to the HFM, that assigns a probability
@@ 1.2 The Hierarchical Feature Model
Marginalising over the low order features $\mathbf{s}_{1:k}$ returns a mixture between a HFM over the remaining $n-k$ features and a frozen state

### v4

@@ A.1 The relevance
This Section recalls the definition of the relevance. We refer to Ref. [9] for a more extended treatment. We describe representations $p(\mathbf{s})$ in terms of the variable $E_{\mathbf{s}}=-\log_{2}p(\mathbf{s})$, which is the minimal number of bits needed to represent state $\mathbf{s}$. The average coding cost $H[\mathbf{s}]=\mathbb{E}\left[E_{\mathbf{s}}\right]$ is the usual Shannon entropy and counts the number of bits needed to describe one point of the dataset. Following Ref. [9], we shall call $H[\mathbf{s}]$ the resolution.
@@
where $p(E)=W(E)2^{-E}$ is the probability that a state $\mathbf{s}$ randomly drawn from $p(\mathbf{s})$ has $E_{\mathbf{s}}=E$, and $W(E)$ is the number of states $\mathbf{s}$ with $E_{\mathbf{s}}=E$. The principle of maximal relevance [9] postulates that maximally informative representations should achieve a maximal value of $H[E]$, which correspond to a uniform distribution of coding costs ($p(E)={\rm const}$ or $W(E)=W_{0}2^{E}$). Representations where coding costs are distributed uniformly should be promoted for the reason that, in an optimal representation, the number $W(E)$ of states $\mathbf{s}$ that require $E$ bits to be represented should match as closely as possible the number ($2^{E}$) of codewords that can be described by $E$ bits. Representation of maximal relevance with a given resolution also have an exponential degeneracy of states $W(E)=W_{0}e^{\nu E}$, with $\nu$ that depends on $H[\mathbf{s}]$. Note that states $\mathbf{s}$ and $\mathbf{s}^{\prime}$ with very different coding costs $E_{\mathbf{s}}$ and $E_{\mathbf{s}^{\prime}}$ can be distinguished by their statistics, because they would naturally belong to different typical sets [^22] By the law of large numbers, typical samples of weakly interacting variables all have approximately the same coding cost, a fact knowns as the asymptotic equipartition property [56]. As in Ref. [66], we take the view that a trained neural network distinguishes the points in a dataset in different typical sets.. Representations that maximize the relevance harvest this benefit in discrimination ability that is accorded by statistics alone.
A.2 The Hierarchical Feature Model
The HFM was introduced in [6], to which we refer for a more complete treatment. The HFM encodes the principle of maximal relevance. It describes the distribution of a string $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ of binary variables that we take as indicators of whether each of $n$ features is present ($s_{i}=1$) or not ($s_{i}=0$). Features are organised in a hierarchical scale of detail and we require that the occurrence of a feature $s_{k}=1$ at level $k$ does not provide any information on whether lower order features are present or not. This means that, conditional on $s_{k}=1$, all lower order features are as random as possible, i.e. $H[\mathbf{s}_{1:k-1}|s_{k}=1]=k-1$ in bits. This requirement implies that the Hamiltonian $E_{\mathbf{s}}$ should be a function of $m_{\mathbf{s}}=\max\{k:~s_{k}=1\}$, with $m_{\mathbf{s}}=0$ if $\mathbf{s}=\mathbf{0}$ is the featureless state [^23] This is because $p(\mathbf{s}|m_{\mathbf{s}}=k)=2^{-k+1}$ so that $p(\mathbf{s})=p(\mathbf{s}|m_{\mathbf{s}})p(m_{\mathbf{s}})=p(m_{\mathbf{s}})/2^{m_{\mathbf{s}}-1}$. ($s_{i}=0~\forall i$). Since there are $2^{k-1}$ states with $m_{\mathbf{s}}=k$, the principle of maximal relevance (i.e. the requirement that $W(E)=W_{0}e^{\nu E}$) excludes all functional forms between $E_{\mathbf{s}}$ and $m_{\mathbf{s}}$ that are not linear. This leads to the HFM, that assigns a probability
@@ A.2 The Hierarchical Feature Model
Marginalising over the low order features $\mathbf{s}_{1:k}$ returns a mixture between a HFM over the remaining nkn-k features and a frozen state